# How to Leverage Predictive Uncertainty Estimates for Reducing Catastrophic Forgetting in Online Continual Learning

### v2

@@ Abstract.
Many real-world applications require machine-learning models to be able to deal with non-stationary data distributions and thus learn autonomously over an extended period of time, often in an online setting. One of the main challenges in this scenario is the so-called catastrophic forgetting (CF) for which the learning model tends to focus on the most recent tasks while experiencing predictive degradation on older ones. In the online setting, the most effective solutions employ a fixed-size memory buffer to store old samples used for replay when training on new tasks. Many approaches have been presented to tackle this problem. However, it is not clear how predictive uncertainty information for memory management can be leveraged in the most effective manner and conflicting strategies are proposed to populate the memory. Are the easiest-to-forget or the easiest-to-remember samples more effective in combating CF? Starting from the intuition that predictive uncertainty provides an idea of the samples’ location in the decision space, this work presents an in-depth analysis of different uncertainty estimates and strategies for populating the memory. The investigation provides a better understanding of the characteristics data points should have for alleviating CF. Then, we propose an alternative method for estimating predictive uncertainty via the generalised variance induced by the negative log-likelihood. Finally, we demonstrate that the use of predictive uncertainty measures helps in reducing CF in different settings.
Uncertainty, Online Continual Learning, Catastrophic Forgetting
conference: 3rd Workshop on Uncertainty Reasoning and Quantification in Decision Making; August 26, 2024; Barcelona, Spain
1. Introduction
Typical machine learning models assume to work in a single-task static scenario where multiple epochs are performed over the same data until convergence. In many real-world situations, however, this setting is too limiting. As an example, let us consider the problem of classifying new COVID-19 variants. Given the evolving nature of the virus, new variants (i.e., new classes) arise over time. In this context, a typical learning model would fail since the standard setting does not consider the dynamic increment of new classes. For this reason, online Continual Learning (online-CL) has been constantly more explored. In online-CL, a single model is required to learn continuously from a sequence of tasks that comes as a stream of tiny batches which can be processed only once (Aljundi et al., 2019, 2017; Mai et al., 2022). Given the overlap between old and new information, the model tends to forget about the past knowledge to focus more on the newest tasks, leading to a performance degradation on previous tasks. This challenge is usually referred to as catastrophic forgetting (CF) (McCloskey and Cohen, 1989; Ratcliff, 1990). Many approaches have been developed to prevent catastrophic forgetting and the most successful ones employ a memory buffer for rehearsal (Chaudhry et al., 2019; Soutif-Cormerais et al., 2023). Memory-based strategies tackle CF by training the model on both current samples and some old samples stored in a limited size memory. In this context, what differentiates each memory-based approach are the retrieval strategy and the update strategy (Mai et al., 2022) — i.e., how to populate and update the memory with meaningful and representative samples, and how to efficiently sample from the memory respectively. Although many approaches have been developed in this direction – ranging from using random approaches (Chaudhry et al., 2019) to exploiting the gradient (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2018b; Aljundi et al., 2019) or the loss (Belilovsky et al., 2019) information – the effect of predictive uncertainty estimates for memory management is largely unexplored. One of the few methods in this direction is presented in Bang et al. (2021), where the authors propose a combination of predictive uncertainty and data augmentation at the memory-level to generate a memory of diverse samples – from the most uncertain to the most representative ones – for each class. Despite remarkable results, the contribution of the uncertainty score on the overall performance is not easily quantifiable. Furthermore, considering the quickly expanding landscape of recently rehearsal-based methods proposed, it is difficult to identify a clear strategy to exploit the memory at its best. In fact, contrasting strategies can be found in the literature. Some of them suggest to use the most representative samples (Yoon et al., 2021; Hurtado et al., 2023), while others consider the samples near the decision boundary as the most useful to reduce CF (Kumari et al., 2022). Hence, it is unclear which type of data points would reduce CF in a consistent manner: Are the easiest-to-forget or the easiest-to-remember samples more effective in combating CF? To answer these and other questions, this paper investigates the contribution of predictive uncertainty scores on reducing CF. Intuitively, using predictive uncertainty to populate the memory represents a solution for identifying the location of the instances in the decision space and considering either the most representative – i.e., the k samples with the lowest uncertainty (bottom-k) – or the marginal ones (i.e., the top-k with the highest uncertainty). In the first part of this work, we evaluate and compare different combinations of uncertainty scores and sorting on CIFAR-10 (Krizhevsky et al., 2009), a dataset commonly used in CL. The objective is to understand how different uncertainty estimates behave when used in different ways under same conditions. In order to achieve the intended goal, following a recent trend on research transparency and comparability (Mundt et al., 2021), the evaluation framework will be freed from any other ’trick’ that could create ambiguity in the assessment. The investigation will focus on describing the characteristics of the considered uncertainty scores, while providing an in-depth analysis of the effect of the metrics under a memory-based regime. In addition to the comparison of established uncertainty scores, we propose a new selection strategy based on recent advances in uncertainty quantification via the logit space. More specifically, we introduce a memory management strategy based on the Bregman Information (BI) as a generalised variance measure, which stems directly from a bias-variance decomposition of the model loss (Gruber and Buettner, 2023).
@@ 1. Introduction
In the second part of our work, we evaluate our findings on more challenging and realistic scenarios. In realistic setups, recent tasks have less data points available than older tasks as the time to collect instances is considerably shorter than for previous tasks. Considering this behaviour, we will focus on long-tailed (LT) datasets with this characteristic. We test our findings on two datasets with controllable degrees of data imbalance, as well as a real-world imbalanced dataset for classification of biomedical images (Yang et al., 2023).
The main contributions of this paper are the followings:
  • We systematically evaluate the performance of established predictive uncertainty scores in reducing CF in online-CL for image classification.
  • We propose an alternative estimate for predictive uncertainty via the generalised variance induced by the loss function.
@@
  • We demonstrate that choosing the most representative samples (bottom-k) to populate the memory consistently reduces CF compared to choosing marginal samples (top-k) – see Figure 1.
  • We validate our findings in realistic scenarios with data imbalance, such as medical image classification.
@@
Figure 1. Relative CF improvement of different uncertainty scores over ER when using the easiest-to-forget (top-k) and the easiest-to-remember (bottom-k) samples.
2. Related Work
We can identify three different scenarios of incremental learning (IL) (van de Ven et al., 2022); task-, domain-, and class-IL. In task-IL, task IDs are explicitly provided along with the data during training and testing. In domain-IL, the problem setting remains unchanged over time but the input distribution (or domain) shifts. Finally, in the class-IL scenario, the model learns to discriminate across a growing number of classes over time. An additional categorization of CL problems considers whether the data can be stored and seen multiple times (offline) or the data are considered as a stream of data which can only be used once (online) (Mai et al., 2022). In this work, we focus on the most challenging and realistic scenario, namely, online class-IL.
In online-CL, where data arrive in single mini-batches and the model is updated with high frequency, memory-based methods are preferred to more sophisticated solutions, e.g., generative methods, for their flexibility and reduced training time (Mai et al., 2022). In this context, the main question is how to optimally manage the memory. In order to reduce catastrophic forgetting, samples in the memory should be representative of their own class, discriminative towards the other classes, and informative enough for the model to recall the information about the old classes. In the literature, many conflicting strategies can be found. Some of them suggest to use the most representative samples (Yoon et al., 2021; Hurtado et al., 2023), while others consider the samples near the decision boundary as the most useful to reduce CF (Kumari et al., 2022).
As already anticipated, Rainbow Memory (RM) (Bang et al., 2021) is the only approach that makes explicit use of uncertainty measures to evaluate the relative position of perturbed samples in the decision space. Given a list $P$ of perturbations, the authors propose to use an agreement score $u_{RM}(x)$ such that, if all the perturbed versions of $x$ are predicted with the same label, then $u_{RM}(x)$ will be 0. Otherwise, the higher is $u_{RM}(x)$ the most uncertain is the model about the considered sample.
@@ 2. Related Work
Noticeable, the definition of online setting in their paper differs from the popular one. In their case, online means that the training stream is only processed once and the memory is updated at the end of each task. In the standard scenario, instead, online implies that only a mini-batch of data points is available to be processed (Mai et al., 2022), and the model and the memory are updated with high frequency (Soutif-Cormerais et al., 2023). In addition to this, to further enhance diversity, they employ data augmentation on the memory. Thus, the contribution of $u_{RM}$ on reducing CF is unclear.
@@
3. Preliminaries and Notation
Following the notation proposed in Bang et al. (2021), we assume to have a set $\bm{\mathcal{C}}=\left\{c_{1},\dots,c_{n}\right\}$ of n different classes. Each class can be randomly assigned to a task $t$. $T_{t}$ represents a subset of classes determined by an assign function $\psi(c)$ such that $T_{t}=\{c|\psi(c)=t\}$. For each task $t$, we have an associated dataset $D_{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ with $x_{i}$ an input sample, $y_{i}$ the corresponding class label, and $n_{t}$ the number of training samples. In the proposed online setting, we assume that samples for each task $t$ come gradually in a stream of mini-batches $b_{t}=\{{(x_{i},y_{i})\in D_{t}\}_{i=1}^{bs}}$ which can be only processed once. It is important to notice that, for each task $t$, $n_{t}$ can vary depending on whether we are working on a class-balanced setting or not. Finally, for replay, we introduce a fixed-size memory buffer $\mathcal{M}$ to store a portion of samples from past tasks. The memory is updated with high-frequency whenever a mini-batch from the stream of data is processed. Following the standard procedure in online-CL, for replay, we assume to extract from the memory a number of samples equal in size to the batch size.
4. Methodology

### v3

@@ Abstract
Many real-world applications require machine-learning models to be able to deal with non-stationary data distributions and thus learn autonomously over an extended period of time, often in an online setting. One of the main challenges in this scenario is the so-called catastrophic forgetting (CF) for which the learning model tends to focus on the most recent tasks while experiencing predictive degradation on older ones. In the online setting, the most effective solutions employ a fixed-size memory buffer to store old samples used for replay when training on new tasks. Many approaches have been presented to tackle this problem and conflicting strategies are proposed to populate the memory. Are the easiest-to-forget or the easiest-to-remember samples more effective in combating CF? Furthermore, it is not clear how predictive uncertainty information for memory management can be leveraged in the most effective manner. Starting from the intuition that predictive uncertainty provides an idea of the samples’ location in the decision space, this work presents an in-depth analysis of different uncertainty estimates and strategies for populating the memory. The investigation provides a better understanding of the characteristics data points should have for alleviating CF. Then, we propose an alternative method for estimating predictive uncertainty via the generalised variance induced by the negative log-likelihood. Finally, we demonstrate that the use of predictive uncertainty measures helps in reducing CF in different settings.
11footnotetext: partner site Frankfurt, a partnership between DKFZ and UCT Frankfurt-Marburg
1 Introduction
Typical machine learning models assume to work in a single-task static scenario where multiple epochs are performed over the same data until convergence. In many real-world situations, however, this setting is too limiting. As an example, let us consider the problem of product recommendation. Since trends may change and new types of products may arrive, new product categories need to be classified. In this context, a typical learning model would fail because the standard setup does not account for the continuous addition of new classes. For this reason, online Continual Learning (online-CL) has been constantly more explored. In online-CL, a single model is required to learn continuously from a sequence of tasks that comes as a stream of tiny batches which can be processed only once (Aljundi et al., 2019b; 2017; Mai et al., 2022). This is to reflect realistic conditions where, for example, new personal data arrive with high-frequency to limited-resources devices (e.g., wearable smart devices) and the model needs to be updated with the incoming data to adjust the model on the fly. In this context, given the overlap between old and new information, the model tends to forget about the past knowledge (which still needs to be classified) to focus more on the newest tasks, leading to a performance degradation on previous tasks. This challenge is usually referred to as catastrophic forgetting (CF) (McCloskey & Cohen, 1989; Ratcliff, 1990).
Many approaches have been developed in online-CL to prevent catastrophic forgetting and the most successful ones fall in the memory-based category. These approaches employ a memory buffer (Chaudhry et al., 2019; Soutif-Cormerais et al., 2023) to a) store samples of past tasks and b) tackle CF by training the model on both current samples and some old samples stored in such limited-size memory (replay or rehearsal). In this context, what differentiates each memory-based approach are the update (or population) strategy and the retrieval (or sampling) strategy (Mai et al., 2022) — i.e., how to populate and update the memory with meaningful and representative samples, and how to efficiently sample from the memory respectively.
Although many approaches have been developed in this direction – ranging from using random approaches (Chaudhry et al., 2019) to exploiting the gradient (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2018b; Aljundi et al., 2019b) or the loss (Belilovsky et al., 2019) information – it is difficult to identify a clear strategy to exploit the memory at its best. In fact, contrasting strategies can be found in the literature. For instance, Kumari et al. (2022) suggest working on instances close to the decision boundary, thus considering the marginal samples as the most important for alleviating CF. Contrarily, in Hurtado et al. (2023) and Yoon et al. (2021), the proposed strategies focus on populating the memory with the most representative samples for each class. Hence, it is unclear which type of data points would reduce CF in a consistent manner: Are the least or the most representative samples more effective in combating CF?
Starting from this question, our investigation seeks to understand the contribution of the most or least representative samples in combating CF through an uncertainty lens. In detail, we will focus on the uncertainty of the model in its predictions, i.e., predictive uncertainty. Intuitively, using measures of uncertainty to populate the memory represents a solution for identifying the location of the instances in the decision space. However, predictive uncertainty can be seen as a composition of aleatoric (data-inherent and irreducible) and epistemic (model-centric and reducible by gathering more data) uncertainty (Malinin & Gales, 2018; Mucsányi et al., 2024), and different uncertainty scores focus more on one source or another leading to different results accordingly. The most common confidence scores mostly capture the (irreducible) aleatoric uncertainty (Wimmer et al., 2023) which is inherent in the data and cannot be reduced by gathering more data. In this case, samples with high confidence from an aleatoric perspective reflect data points that are distant from the decision boundary irrespective of the fact that they may be outliers or not. In the epistemic case, instead, samples with high confidence (i.e., low epistemic uncertainty) indicate that the observed data sufficiently support the inferred patterns (or, in other words, that the samples are representative of the data distribution). Thus, although representativeness is an intrinsic property of the data, we can interpret (low) epistemic uncertainty as a way to infer this property from a predictive model (see Section 4.2.3 for further details). For this reason, we hypothesize it would be beneficial to focus on samples with a low epistemic uncertainty. To empirically validate this hypothesis, we propose a new population strategy based on recent theoretical contributions in uncertainty quantification via the Bregman decomposition. More specifically, we introduce a memory management strategy based on the Bregman Information (BI) as a generalised variance measure, which stems directly from a bias-variance decomposition of the model loss (Gruber & Buettner, 2023). Based on this insight, we interpret BI as a measure of epistemic uncertainty that is statistically well grounded in a bias-variance decomposition, and propose to use it for populating the memory.
@@ 1 Introduction
In the first part of this work, we evaluate and compare different combinations of uncertainty scores and sorting on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), two datasets commonly used in CL. The objective is to understand how different uncertainty estimates behave when used in different ways under same conditions. In order to achieve the intended goal, following a recent trend on research transparency and comparability (Mundt et al., 2021), the evaluation framework will be freed from any other ’trick’ that could create ambiguity in the assessment. The investigation will focus on describing the characteristics of the considered uncertainty scores, while providing an in-depth analysis of the effect of the metrics under a memory-based regime. In the second part of our work, we evaluate our findings on more challenging and realistic scenarios. In realistic setups, recent tasks have less data points available than older tasks as the time to collect instances is considerably shorter than for previous tasks. Considering this behaviour, we will focus on long-tailed (LT) datasets with this characteristic. We test our findings on two datasets with controllable degrees of data imbalance, as well as a real-world imbalanced dataset for classification of biomedical images (Yang et al., 2023). The goal of this work is not to propose a new method that outperforms the state-of-the-art, but rather to present practical insights from an uncertainty-aware perspective on the desirable characteristics samples should have to alleviate catastrophic forgetting in both standard and realistic scenarios.
Thus, the main contributions of this paper can be summarised as follows:
  • We conduct a systematic evaluation of established predictive uncertainty scores to assess their effectiveness in mitigating catastrophic forgetting (CF) in online continual learning for image classification tasks.
  • We propose the use of Bregman Information (BI) as measure of epistemic uncertainty to populate the memory in online-CL.
@@
  • As summarised in Figure 1, we demonstrate that a) selecting the most representative samples (bottom-k) to populate the memory consistently reduces CF, as compared to choosing marginal samples (top-k); and b) focusing on estimates of epistemic uncertainty (BI) provides an advantage compared to aleatoric uncertainty estimates.
  • Different from conventional evaluation pipelines, we validate our findings on both standard benchmarks and realistic scenarios, including challenging settings like medical image classification with imbalanced data.
@@
Figure 1: Relative CF improvement of different uncertainty scores over ER when using the easiest-to-forget (top-k) and the easiest-to-remember (bottom-k) samples. The bottom-k strategy (in darker colour) reduces CF in all cases. Furthermore, the proposed BI-based uncertainty estimate (in green) further diminishes CF compared to common uncertainty scores.
2 Related Work
2.1 Continual Learning
According to van de Ven et al. (2022), three different scenarios of incremental learning (IL) can be identified; 1) Task-IL where the task ID information is available at both training and testing time; 2) Domain-IL in which the learning task remains unchanged (e.g., binary classification) but there is a shift in the input distribution; and 3) Class-IL where the number of classes to discriminate can grow over time. Additionally, CL problems can be further categorised depending on whether the data can be accessed and processed multiple times (offline) or a single-pass through the data is expected (online) (Mai et al., 2022). In this work, following a popular trend in the literature, we focus on the most challenging scenario, i.e., online class-IL.
2.2 Class-Incremental Learning
Following the classification proposed by Mai et al. (2022), class-IL approaches can be grouped into the following categories: a) Regularization techniques that adjust the model parameter updates by incorporating penalty terms in the loss function (Aljundi et al., 2018; Lee et al., 2017), modifying parameter gradients during optimization (Chaudhry et al., 2018b; He & Jaeger, 2018), or employing knowledge distillation (Wu et al., 2019; Rannen et al., 2017); b) Memory-based techniques in which a fixed-size subset of past samples is stored for replay (Aljundi et al., 2019a; Chaudhry et al., 2019) or for regularization purposes (Nguyen et al., 2018; Tao et al., 2020) ; c) Generative-based techniques which involve training generative models to produce pseudo-samples that replicate the information from past tasks (Lesort et al., 2019; Shin et al., 2017), and d) Parameter-isolation-based techniques that allocate distinct model parameters to each task, either by activating only the relevant parameters for each task (Fixed Architecture) (Mallya & Lazebnik, 2018; Serra et al., 2018) or by adding new parameters while keeping the existing ones unchanged (Dynamic Architecture) (Yoon et al., 2018; Aljundi et al., 2017).
In online-CL, where data are received in mini-batches and the model is updated with high frequency, rehearsal-based methods are favoured over more complex solutions like generative methods due to their flexibility and reduced training time (Mai et al., 2022).
2.2.1 Population Strategies in Rehearsal-based CL
As already anticipated, there exist methods that use different metrics for deciding the samples to store in the memory. In Aljundi et al. (2019b), the gradient information is used as a feature to maximize the diversity of the samples in the memory. In Chaudhry et al. (2018b), the gradient of the current mini-batch is compared with with the gradient of a randomly sampled set of the same size from the memory. If the dot product between the current gradient and the memory gradient is negative, the current gradient is projected. Otherwise, the gradient is used normally. Similarly, in Yoon et al. (2021), gradient vectors are used to select the most representative and informative coreset at each iteration. Other methods exploit the loss information to select samples based on their interference on the loss function (Aljundi et al., 2019a) or to analyze the loss region for improving generalization and minimizing CF (Verwimp et al., 2021). Sun et al. (2021) and Wiewel & Yang (2021) address the problem of memory diversity from an information-based perspective via entropy-based functions. As previously mentioned, the characteristics samples should have to be stored in the memory and alleviate CF are not clear. For instance, Hurtado et al. (2023) propose a method to eliminate outliers from the memory such that only the most representative samples for each class are kept via a label-homogeneity score. Intuitively, their approach inspect the neighborhood of a given sample to check whether it is surrounded by samples from different classes or not. If the label homogeneity score is low (i.e., most of the nearest samples do not share the same class label), it means we are in the presence of an outlier or noisy sample and the considered sample should be discarded from memory sampling. Contrarily, in Kumari et al. (2022), the authors claim that the replay-phase should focus on marginal samples and propose a method to synthesize samples near the forgetting boundary that are confused with the current task. The approach consists on identifying the easiest-to-forget samples in the memory and then on moving those samples closer to the data points of the current task. Finally, the new perturbed samples are used for replay to cover marginal samples more consistently.
In both cases, the idea is similar to using uncertainty estimates to detect data points with specific characteristics – i.e., outliers or representative samples. Differently from their approaches, which require the samples’ vector representation, our method exploits predictive uncertainty for the same purpose. It can be seen as a more principled way of identifying whether a sample is representative of its class or not. Nevertheless, despite the use of predictive uncertainty looks appealing, its usage for populating the memory with samples with desired properties is largely unexplored.
2.2.2 Uncertainty-aware Memory Management in CL
The idea of employing predictive uncertainty for memory management is borrowed from Active Learning (AL). In this context, uncertainty sampling (Lewis & Gale, 1994) is used to decide which samples from a pool of unlabeled data will be considered for labeling. The main assumption is that the most uncertain samples represent the most informative data points. Thus, by including them in the training set, we can improve the overall performance. Popular examples of uncertainty scores include, to name a few, entropy, smallest margin, and least confidence (Shannon, 1948; Campbell et al., 2000; Lewis & Gale, 1994; Culotta & McCallum, 2005).
Analogously, we can use estimates of predictive uncertainty for populating memory buffers in replay-based methods for online-CL. In Bang et al. (2021), the authors argue that samples stored in the memory should be representative of their own class, but also discriminative towards the remaining ones. Starting from this assumption, they evaluate the relative position of the samples in the decision space by estimating the predictive uncertainty of perturbed samples; exemplars with high uncertainty should be closer to the decision boundary, while the most certain ones should be located closer to the center of the corresponding class distribution. To ensure diversity, they use a step-sized sampling which guarantees to keep in the memory samples that span from the most uncertain to the most certain ones. Given a list $P$ of perturbations, the authors propose to use the following uncertainty score $u_{RM}(x)$:
$u_{RM}(x)=1-\frac{1}{P}\max_{c}S_{c}$ (1)
where $S_{c}=\sum_{i=1}^{P}\mathbbm{1}_{c}\operatorname*{argmax}_{\hat{c}}p(y=\hat{c}|\tilde{x}_{i})$. Eq. (1) represents an agreement score with respect to the perturbations. If the predicted classes $\hat{c}$ of all the perturbed versions of $x$ (i.e., $\tilde{x}_{i}$) are predicted with the same label $c$, then $u_{RM}(x)$ will be 0. Otherwise, the higher is $u_{RM}(x)$ the most uncertain is the model about the considered sample.
@@ 2.2.2 Uncertainty-aware Memory Management in CL
The authors claim to work in the online setting. Nevertheless, the definition of online setting in the paper differs from the popular one. In their case, online means that the training stream is only processed once and the memory is updated at the end of each task. In the standard scenario, instead, online implies that only a mini-batch of data points is available to be processed (Mai et al., 2022), and the model and the memory are updated with high frequency (Soutif-Cormerais et al., 2023). In addition to this, to further enhance diversity, they employ data augmentation on the memory. Thus, the contribution of Eq. (1) on reducing CF is unclear.
@@
3 Preliminaries and Notation
Following the notation proposed in Bang et al. (2021), we assume to have a set $\bm{\mathcal{C}}=\left\{c_{1},\dots,c_{n}\right\}$ of n different classes. Each class can be randomly assigned to a task $t$. $T_{t}$ represents a subset of classes determined by an assign function $\psi(c)$ such that $T_{t}=\{c|\psi(c)=t\}$. For each task $t$, we have an associated dataset $D_{t}=\{(x_{i},y_{i})\}_{i=1}^{n_{t}}$ with $x_{i}$ an input sample, $y_{i}$ the corresponding class label, and $n_{t}$ the number of training samples. In the proposed online setting, we assume that samples for each task $t$ arrive sequentially in a stream of mini-batches $b_{t}=\{{(x_{i},y_{i})\in D_{t}\}_{i=1}^{bs}}$, with each mini-batch available for a single pass. It is important to notice that, for each task $t$, $n_{t}$ can vary depending on whether we are working on a class-balanced setting or not. Finally, for replay, we introduce a fixed-size memory buffer $\mathcal{M}$ to store a portion of samples from past tasks. The memory is updated with high-frequency whenever a mini-batch from the stream of data is processed. Following the standard procedure in online-CL, for replay, we assume to extract from the memory a number of samples equal in size to the batch size.
4 Methodology